Steam at 100°C is discharged into a container filled with water at 50°C. The initial volume of water is 2000 cm3. After a certain period of time, it is observed that the volume of water has increased, and its final temperature is 70°C.
(a)Why does this happen?
(b)Calculate the increased volume of water.
1
Expert's answer
2020-12-06T17:21:33-0500

a) Some amount of steam is condensed onto water.


b)


"mR=mc(T_s-T_f)+m_wc(T_f-T_w)\\\rho VR=\rho Vc(T_s-T_f)+\rho V_wc(T_f-T_w)\\
 VR= (V(T_s-T_f)+V_w(T_f-T_w))c\\(2260000)V= 4200(100-70)V\\+2000(70-50)(4200)\\V=78.7\ cm^3"